What's the FCFF of a company with total revenues of $857 million, operating profit margin of 30%, tax rate of 32% and reinvestment rate of 33%? Answer in millions, rounded to one decimal place.
Answer #1

Future cash flow(=EBIT*(1-T))-Reinvestment

Total revenue=857mn,

Operating profit margin=operating profit/revenue

0.3=Operating profit/857

EBIT=0.3*857=257.1

=EBIT*(1-T))-Reinvestment rate

=2571*(1-0.32)=174.828

=174.828-(174.828*0.33)

=117.135.

117.1 million.